OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -6) and a directrix at y = 6.

OpenStudy (anonymous):

you know what this looks like? if you do, then we can do it quickly

OpenStudy (anonymous):

no

OpenStudy (anonymous):

would it just be a x^2 parabola but the starting point is shifted 6 to the down?

OpenStudy (anonymous):

i meant down

OpenStudy (anonymous):

no

OpenStudy (anonymous):

then i dont know

OpenStudy (anonymous):

lets draw a picture with the focus and the directrix only

OpenStudy (anonymous):



OpenStudy (anonymous):

so it isnt a parabola?

OpenStudy (anonymous):

yes, it is the vertex is half way between the focus and the directrix. so the vertex is the origin \((0,0)\)

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

but you should see from the picture that the parabola opens down, not up

OpenStudy (anonymous):

since -6 that means opens down?

OpenStudy (anonymous):



OpenStudy (anonymous):

ok i see

OpenStudy (anonymous):

but that does not make is \(y=-x^2\) we still have to be careful

OpenStudy (anonymous):

y=-(x-6)^2?

OpenStudy (anonymous):

general form is \[4p(y-k)=(x-h)^2\] in your case \((h,k)\) is \((0,0)\) so we are at \[4py=x^2\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

your answer would shift \(y=-x^2\) to the right 6 units, but that is not what we are trying to do all you need now to finish \[4py=x^2\] is find \(p\) which is the distance between the vertex and the focus that distance is pretty clearly \(6\)

OpenStudy (anonymous):

so it would be y=x^2/24?

OpenStudy (anonymous):

don't forget the minus sign

OpenStudy (anonymous):

yeah i forgot from the 24

OpenStudy (anonymous):

\[-24y=x^2\] or \[y=-\frac{x^2}{24}\]

OpenStudy (anonymous):

want to check it is right?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

i wrote "parabola" and then the equation it gives the focus, directrix, vertex, etc http://www.wolframalpha.com/input/?i=parabola+-24y%3Dx^2

OpenStudy (anonymous):

ohh wow thats convenient lol